OpenStudy (anonymous):

A cola company wants to test their cola against their biggest competitor's cola. They decide to randomly sample 150 people. The results from their survey show that 98 people out of 150 people prefer their cola to their competitor's. What is the best prediction of the number of people who will prefer their cola if 1,200 people are asked? 784 1,078 980 686

OpenStudy (anonymous):

i tho 686

OpenStudy (amistre64):

what are your thoughts?

OpenStudy (amistre64):

id set up a proportion ...

OpenStudy (anonymous):

u put it in a fraction

OpenStudy (amistre64):

yep if 98 out of 150, then x out of 1200 98/150 = x/1200

OpenStudy (anonymous):

my problem is not good at multiplying

OpenStudy (amistre64):

cant help too much with that .. how many times does 150 go into 1200? 1200/150 120/15, 40/5 = 8 so 8 times as many people ... 98*8 is? well, 9*8 is 72 so its a little bigger than 720

OpenStudy (amistre64):

of since 98 is about 100, is a little less then 8*100 ...

OpenStudy (anonymous):

thanks thank got it in new it was a fraction one